Mark has a choice to play a game and pay $80 with a probability of 40% to win $250. Should he play or not? Support your answer by constructing a decision tree with EMV calculations for decision options.
Answer #1

P(Win) = 0.40

If Mark wins, he earns = 250 - 80 = 170

P(Loss) = 1 - P(Win) = 0.60

If Mark losses, he earns = $ -80

EMV = 0.40*170 + (-80)*0.60 = 68 -48

EMV = $ 20

Since EMV is positive so Mark should play the game